OpenStudy (anonymous):

Arrange these numbers from least to greatest:

OpenStudy (anonymous):

\[\frac{ 3 }{ 5 }\]\[\frac{ 5 }{ 7 }\]\[\frac{ 2 }{ 9 }\]

OpenStudy (anonymous):

first make sure the denominators are the same .

OpenStudy (hyperdog3):

3/5 = 0.60 5/7 = 0.71 2/9 = 0.20 So first divide them and get decimals so it's easier and now you can order them from least to greatest. 2/9, 3/5, 5/7

OpenStudy (anonymous):

Thanks. @igotoschool && @hyperdog3